• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eric Voutier

    Log entry time 21:49:33 on June14,2005

    Entry number 146161

    keyword=Shift Summary


    New crew on shift: Vincent Sulkowsky, Drew Whitbeck, Eric Voutier
    -----------------

    16:00 Coming on shift while magnet cycling is underway

    16:15 Q3 left arm is down, power off going to zero.

    16:30 Q3 reset does not work; going to control access for hard reset

    16:45 MCC asks to delay the access for 20mn to finish some work on
    the energy spread

    17:25 Going down to the hall to reset Q3; following the reset
    procedure downstairs did not succeed; the faults came back every
    time one tries to put the power supply on.

    18:15 Heidi is down in the hall trying to solve the problem

    18:27 The left arm is ramping down... Q3 problem is serious, needs
    about 12 hours to fixe it when the parts are on site (resistors in
    the breaker).

    20:30 Going to Rob plan (entry 146136)

    20:50 Beam is back, following the plan

    21:37 Parking the RICH, HV off, gas flow @ 410
    Going to empty target

    21:50 ... Stopping the shift, Eugene should come over the night to
    do some testing

    Run History
    -----------

    I = 50uA

    Run Number | Tachometer speed | T1 rate
    parity/coda | |
    -----------------------------------------
    3437/4596 ....... 750 .......... 153.4 kHz
    3438/4597 ....... 600 .......... 156.6 kHz
    3439/4598 ....... 500 .......... 125.6 kHz
    3440/4599 ....... 400 .......... 81.1 kHz
    3441/4600 ....... 300 .......... 82.8 kHz